A helicopter pilot throws a ball downwards from an unknown height above the ground. The ball has an initial speed of 14.1 m/s. The ball hits the ground 9.2 s later. Find the initial height of the ball above the ground


1
Expert's answer
2022-04-20T13:46:59-0400

"Let~v_0=14.1~m/s,~t=9.2~s,~g=9.81~m/s^2\\
h=v_0t+gt^2/2=(v_0+gt/2)t=\\
=(14.1+9.81\cdot9.2/2)\cdot9.2~(m)\approx545~(m)"